# SALAD: Achieve High-Sparsity Attention via Efficient Linear Attention Tuning for Video Diffusion Transformer

### v1

@@ 5 Experiments
Table 1: Quality and efficiency result of SALAD and baselines.
Method SC (↑) BC (↑) IQ (↑) TC (↑) VisionReward (↑) Sparsity (↑) Speedup (↑)
Original (Full Attention) 95.88 96.17 65.93 25.31 0.109 0%
Training-free
[…]
Tuning-based
@@
5.1 Setup
Base Model & Datasets. We test our method and along with other baseline methods, using the Wan2.1-1.3B[20] model with a generation configuration of 480p resolution and 77 frames. For our method and training-based baseline methods, we use 2000 videos from Mixkit, an open-source video dataset released with Opensora plan [11], for model fine-tuning with 1600 steps.
Baselines. We compare our method with various efficient video generation approaches. For training-free baselines, we include two static sparse attention methods: spatio-temporal sliding window attention and PAROAttention [40]—as well as one adaptive sparse attention method, SVG2 [28]. The spatio-temporal local attention is extended from DiTFastAttnV2, incorporating the token reordering kernel from SVG. Following common practice, all training-free methods retain full attention for the denoising steps close to the noise to preserve generation quality. For tuning-based baselines, we evaluate ST-SWA tuned with LoRA, SLA [34] and Top-K sparse attention implemented with VMoBA [23] mentioned in Sec. 3.1. Both methods are fine-tuned under the same computational budget and follow their original training configurations to ensure a fair comparison. For end-to-end speedup evaluation, we adopt the official open-source kernels released by these methods.
Evaluation Metrics. We evaluate the generated video quality using the VBench [8] and Vision Reward [27] metrics. From VBench, we report Subject Consistency (SC), Background Consistency (BC), Image Quality (IQ), and Text Consistency (TC; also reported as Overall Consistency in other works). We follow the official VBench settings and use its extended prompt set to generate videos. Comparisons with baseline video quality and the initialization study are conducted on the full set of VBench prompts. Besides, parameter-efficiency comparisons with LoRA and other ablation studies are performed on a subset of VBench prompts. For efficiency evaluation, we report Sparsity (the proportion of computations reduced by sparse attention) and end-to-end speedup. All speedup measurements are carried out on a single GPU with a batch size of 1.
5.2 Main Results
Comparison with Sparse Attention Methods. We compare SALAD with state-of-the-art sparse attention methods, as summarized in Tab. 1. When applied to the spatial-temporal model, SALAD consistently outperforms prior methods across all four evaluation metrics, achieving a $1.77\times$ speedup. Existing post-training sparse attention approaches with roughly 50% sparsity suffer from noticeable quality degradation and limited speedup. Although SLA [34] effectively reduces training cost, it still requires 20,000 video samples and 2,000 training steps with a batch size of 64 to maintain generation quality, which leads to performance degradation under our training budget. LoRA tuning recovers image quality and text consistency, but subject and background consistency remain below the dense baseline at 63% sparsity. In contrast, SALAD achieves 90% sparsity while maintaining video quality comparable to, and even surpassing the dense model, achieving 96.01 in subject consistency and 25.48 in text consistency.
Method Subject Background Image Text Trainable
Consistency Consistency Quality Consistency Parameters
Original (Full Attention) 94.19 96.76 70.17 28.95 -
@@ 5.2 Main Results
Comparison with LoRA. We compared SALAD with various LoRA configurations under the same sparsity. As shown in Tab. 7, increasing the number of LoRA parameters during model tuning does not consistently improve video quality, with the best performance achieved at rank 128. SALAD outperforms almost all the LoRA variants in Subject Consistency, Background Consistency, and Image Quality, while maintaining comparable Text Consistency, using fewer trainable parameters than LoRA at rank 256. This suggests that the observed gains in video quality arise from the method’s intrinsic design rather than an increase in parameter count.
@@ 5.3 Ablation Study
Architecture Design Tab. 3 presents an ablation study on how each design component affects overall video quality. Adding the projection alone improves subject consistency and image quality but reduces background and text consistency. Incorporating the gate alongside the projection enhances performance across all metrics, indicating that the projection alone is insufficient to balance branch outputs. Compared with the Non-shared weight variant, our design achieves even higher scores on all metrics while introducing substantially fewer parameters, demonstrating that the combination of shared architecture, projection, and gating not only optimizes performance but also provides a favorable trade-off between accuracy and parameter efficiency.
Table 3: Performance comparison across different architecture designs. Here, $D$ denotes the number of channels of the input hidden states to the attention block, and $H$ denotes the number of channels used for the query/key/value. In our deployment, we set $D=H$. For the parameter count of each projection matrix, $W^{Q,K,V,O}$ is $H\times D$, the projection layer for the linear attention branch introduces $H\times D$ parameters, and the weight of linear layer in GATE operation has $D\times 1$ parameters.
@@
Shared + Proj 95.94 95.22 68.72 23.14 $H\times D$
@@
Initalization. We experiment SALAD with different initialization settings. Zero initialization continuously has lower training loss compared with random initialization. After finetuning, the model with zero initialization projection module demonstrates higher subject, background, and text consistency as shown in Tab. 4.
Gate Design. To further investigate how the gate facilitates performance recovery, we examine two variants. The Constant Scale variant replaces the gate with a fixed scalar $\lambda=0.5$, while the Detached Gate variant is designed to analyze whether backpropagating gradients to the input positively contribute to performance. In this variant, such gradients are blocked, while the gate parameters remain trainable. As shown in Tab. 9, the Constant Scale variant improves video quality, indicating that adjusting the contribution of the linear attention during both training and inference enhances recovery. However, it performs worse than the input-dependent gate, suggesting that adaptive gating more effectively regulates the auxiliary role of the linear attention branch. The Detached Gate variant also achieves higher video quality, implying that the improvement arises primarily from dynamic scaling of the linear attention rather than from gradient flow to the input.
Table 4: Performance comparison across initialization settings.
Initialization SC (↑) BC (↑) IQ (↑) TC (↑)
Random 96.16 95.51 65.67 25.70
@@
Table 5: Performance comparison across gate design.
Gate Design SC (↑) BC (↑) IQ (↑) TC (↑)
w/o Gate 95.94 95.22 68.72 23.14
[…]
Constant Scale 97.45 96.55 64.96 24.23
@@ 5.4 Additional Observations
Figure 7: Generated Video Samples of LoRA, SALAD, and SALAD with linear branch dropped during inference.
Effect of Branch To investigate the effect of the branch, we dropped the branch in inference after using SALAD (denote as SALAD-drop) and compared the video generated with SALAD and the one using LoRA. Fig. 7 shows that, without the branch, some prompt objects (e.g., fishing boats, top) are missing, and subject duplication occurs (e.g., two dogs, bottom) in the LoRA-tuned video, suggesting that LoRA tuning cannot fully recover the long dependency loss due to the high sparsity. The generated videos of SALAD-drop have better video-text consistency. However, there is a loss of details (e.g., the boat in video sample 1 (upper) and the dog’s tail in video sample 2 (bottom)). With the linear attention branch, the videos show more details and better frame continuity. This demonstrates that adding a linear branch allows the sparse attention component to focus on basic video generation performance restoration, while the linear branch adds additional token interactions to handle detail supplementation and coherence restoration function.
Figure 8: Examples of linear and sparse attention maps at block 8, timestep 20, head 2 and 4. Linear attention attends to long distant tokens, adding global information.
Attention Map Analysis We visualize attention maps from selected heads of sparse and linear attention maps in Fig. 8. In both examples, linear attention weight expands more extensively across longer token distances — an observation evident from the lighter, more distributed regions in its maps. The extended range of linear attention allows it to capture long-range dependencies that are largely neglected by sparse attention, enhancing its capacity for global context integration.
Post-training Branch Dropping We find that the linear attention branch on different attention layers benefits the performance differently, and some linear branches can be dropped while maintaining performance. As shown in the Appendix A.3, dropping 20% of branches with larger values achieves performance comparable to not dropping while providing an additional 5% acceleration. These results suggest that more fine-grained branch-dropping strategies could further improve the efficiency of SALAD, which we leave for future work.
@@ 7 Acknowledge
Appendix A Appendix
A.1 Sparse Attention Implementation Details
A.1.1 Spatial-Temporal Sliding Window Attention
As mentioned in the Sec. 3.1, we now provide more details about the Spatial-Temporal Sliding Window Attention (ST-SWA).
Previous studies [24] have shown that attention heads in Video Diffusion Transformers exhibit two characteristic sparse patterns: spatial local attention, where a token primarily attends to tokens within the same frame, and temporal attention, where a token focuses on tokens at the same spatial location across neighboring frames. In these models, 3D video latents of size $H_{h},W_{h},F_{h}]$ are flattened into a 1D token sequence.
@@ A.1.1 Spatial-Temporal Sliding Window Attention
In our implementation of spatial-temporal attention, we adopt the headwise permutation kernel from SVG [24] and the headwise sliding-window attention kernel from DiTFastAttnV2 [33]. For configuring the window size of each attention head, we follow the procedure of DiTFastAttnV2 and use eight VBench video prompts as the profiling dataset. We then greedily select the smallest window size whose induced error does not exceed a pre-defined threshold. The error threshold $\delta$ (maximum RSE after applying the sparse attention) is set to $2.0$.
A.1.2 Top-K Block Sparse Attention
@@
A.2 Training Details
A.2.1 Training Settings

### v2

@@ 5 Experiments
Table 1: Quality and efficiency results of SALAD and baselines on WAN1.3B. * indicates the sparse attention is applied to a part of diffusion steps. Following the SVG2 [yang2025sparse] practice “with sparse attention skipped during the first 30% of denoising steps for all methods, as these steps are critical for generation quality”, several baselines adopt the same setting.
Tokens Method Category SC↑ BC↑ IQ↑ TC↑ VR↑ Sparsity↑ Speedup↑
30k Original (Full Attention) 95.88 96.17 65.93 25.31 0.109 0%
@@
[…]
100k Original (Full Attention) 92.09 95.47 65.29 27.50 0.112 0%
[…]
5.1 Setup
Base Model & Datasets.
We evaluate our method against representative baselines on Wan2.1-1.3B [wan2025] at 480p with 77 frames (30k tokens). We also extend Wan-1.3B to 720p (100k tokens) with SALAD, and apply our method to Wan2.1-14B at 480p. All training-based methods are fine-tuned on 2,000 Mixkit videos under the OpenSora plan [lin2024open].
Baselines. We compare our method with various efficient video generation approaches. For training-free baselines, we include two static sparse attention methods: spatio-temporal sliding window attention, PAROAttention [zhao2025paroattention] and SVG2 [yang2025sparse]. The spatio-temporal local attention is extended from DiTFastAttnV2, incorporating the token reordering kernel from SVG. Following common practice [xi2025sparse, yang2025sparse], some of training-free methods retain full attention for the denoising steps close to the noise to preserve generation quality. For tuning-based baselines, we evaluate SLA [zhang2025sla], ST-SWA tuned with LoRA and Top-K sparse attention implemented with VMoBA [wu2025vmoba] mentioned in Sec. 3. Both methods are fine-tuned under the same computational budget and follow their original training configurations to ensure a fair comparison. For end-to-end speedup evaluation, we adopt the official open-source kernels released by these methods.
Training Configuration. For all tuning-based methods, we apply LoRA to the attention projection layers ($W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$) with rank $r=128$ and $\alpha=256$. For SALAD, we additionally fine-tune the scale-related modules (see Fig. 3). All models are trained with a batch size of 8. For Wan2.1-1.3B, training takes 1,600 steps ($\sim$20 GPU hours) at 480p and 1,000 steps ($\sim$30 GPU hours) at 720p. For Wan2.1-14B at 480p, training takes 600 steps ($\sim$20 GPU hours).
Evaluation Metrics. We evaluate the quality of generated videos using the VBench [huang2024vbench] and Vision Reward [xu2024visionreward]. From VBench, we report Subject Consistency (SC), Background Consistency (BC), Image Quality (IQ), and Text Consistency (TC). We use the extended prompt set of VBench to generate videos and generate 2 videos for each prompt. Comparisons with baseline video quality on WAN1.3B and the initialization study are conducted on the full set of VBench prompts. Besides, WAN14B evaluation, parameter-efficiency comparisons with LoRA and other ablation studies are performed on a subset of VBench prompts. All ablation studies are conducted on WAN1.3B with 480p resolution. For efficiency evaluation, we report Sparsity (the proportion of computations reduced by sparse attention) and end-to-end speedup. All speedup measurements are carried out on a single GPU with a batch size of 1.
5.2 Main Results
Comparison with Sparse Attention Methods. We compare SALAD with state-of-the-art sparse attention methods at both 30k and 100k token lengths, as summarized in Tab. 1. In the 30k-token setting, when applied to the spatial-temporal SWA, SALAD consistently outperforms prior approaches across all four evaluation metrics, achieving a $1.72\times$ speedup. Existing post-training sparse attention methods exhibit noticeable quality degradation at 90% sparsity, even when retaining full attention for a subset of diffusion steps. Although SLA [zhang2025sla] reduces training cost, it still requires 20,000 video samples and 2,000 training steps with a batch size of 64 to preserve generation quality, and degrades under our limited training budget. LoRA tuning partially restores image fidelity and text alignment; however, the overall performance remains below the dense baseline, even when 90% sparse attention is applied to only 70% of diffusion timesteps (63% overall sparsity). In contrast, SALAD achieves 90% overall sparsity while maintaining video quality comparable to, and even surpassing, the dense model on VBench. In the 100k-token setting, SALAD achieves a $2.03\times$ speedup while preserving performance comparable to full-attention baselines.
Notably, SALAD extends existing sparse attention methods and adopts their released CUDA kernels. Consequently, its acceleration depends on kernel efficiency.The Top-K sparse attn is built upon VMoBA [wu2025vmoba], whose released kernel is not fully optimized, thereby limiting the observed speedup of Top-K.
We also conduct a preliminary experiment on Wan2.1-14B. As shown in Tab. 2, using ST-SWA with SALAD and post-tuning branch dropping achieves about $1.59\times$ speedup while maintaining performance close to the full-attention model, demonstrating the feasibility of SALAD on large-scale models.
@@ 5.2 Main Results
Table 2: Quality and efficiency comparison on WAN14B.
Method SC↑ BC↑ IQ↑ TC↑ VR↑ Sparsity↑ Speedup↑
Original (Full Attention) 94.72 97.53 62.62 27.55 0.0938 0%
[…]
Table 3: Performance and trainable parameters comparison with LoRA.
Method SC (↑) BC (↑) IQ (↑) TC (↑) Trainable Parameters
Original (Full Attention) 94.19 96.76 70.17 28.95
@@
Comparison with LoRA. We compared SALAD with various LoRA configurations under 90% sparsity. As shown in Tab. 9, increasing the number of LoRA parameters during model tuning does not consistently improve video quality, with the best performance achieved at rank 128. SALAD outperforms almost all the LoRA variants in SC, BC, and IQ, while maintaining comparable TC, using fewer trainable parameters than LoRA at rank 256. This suggests that the observed gains in video quality arise from the method’s intrinsic design rather than an increase in parameter count.
@@ 5.3 Ablation Study
Multi-level Dynamic-Static Scaling. Tab. 4 presents an ablation study on how each design component affects the video quality. Adding the layer-wise static scaling projection alone improves subject consistency and image quality but reduces background and text consistency. Incorporating the input-timestep-aware dynamic scaling alongside the projection enhances performance across all metrics, indicating that the dynamic scaling is essential for coordinating the two branches’ output. Compared with the Non-shared weight variant (i.e., sparse and linear attn use independent $W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$), our design achieves higher scores on all metrics while introducing substantially fewer parameters, demonstrating that SALAD not only optimizes performance but also provides a favorable trade-off between accuracy and parameter efficiency.
Table 4: Performance comparison across different architecture designs. Here, $D$ denotes the number of channels of the input hidden states and Q/K/V. For the parameter count of each projection matrix, $W^{Q,K,V,O}$ is $D\times D$, the static scaling projection layer for the linear attention branch introduces $D\times D$ parameters, and the weight of linear layer in dynamic scaling operation has $D\times 1$ parameters.
@@
Shared + Static Scaling 95.94 95.22 68.72 23.14 $D\times D$
@@
Initalization of Static Scaling layers. We experiment SALAD with random initialization settings when applying static scaling. Zero initialization has lower training loss compared with random initialization. After finetuning, the model with zero initialization projection module demonstrates higher subject, background, and text consistency as shown in Tab. 5 (a).
Input-Timestep-Aware Scaling. To further investigate how the dynamic scaling facilitates performance recovery, we examine two variants. Constant Scale variant replaces the dynamic scaler with a fixed scalar $\lambda=0.5$. The Detached variant blocks the gradient flow from the Dynamic Scaler to the input features while keeping the Scaler’s internal channel-mixing linear layers trainable. By truncating gradient flow to the input, this variant helps determine if backpropagating gradients to the backbone provides any additional benefit beyond the dynamic scaling itself.
As shown in Tab. 8 (b), Constant Scale improves video quality, indicating that adjusting the contribution of the linear attention during both training and inference enhances recovery. However, it performs worse than the dynamic scaling values, suggesting that adaptive scaling more effectively regulates the auxiliary role of the linear attention branch. Detached variant also achieves higher video quality, implying that the improvement arises primarily from dynamic scaling of the linear attention rather than from gradient flow to the input.
Table 5: Ablation study on scaling related modules. (↑) indicates higher is better.
@@ 5.4 Additional Observations
Effect of the Branch. To analyze the role of the linear branch, we disable it at inference after SALAD training (denoted as SALAD-drop) and compare the results with full SALAD and LoRA-only tuning. As shown in Fig. 5, LoRA-only tuning fails to recover long-range dependencies under high sparsity, leading to missing prompt objects (e.g., fishing boats, top) and subject duplication (e.g., two dogs, bottom). SALAD-drop improves video–text consistency but exhibits noticeable detail loss (e.g., the boat in sample 1 and the dog’s tail in sample 2). In contrast, the full SALAD model produces richer details and better temporal continuity. These results indicate that the sparse attention primarily restores core generation quality, while the linear branch provides complementary token interactions to enhance fine details and coherence.
@@ 7 Acknowledge
Appendix 0.A Appendix
0.A.1 Sparse Attention Implementation Details
0.A.1.1 Spatial-Temporal Sliding Window Attention
As mentioned in the Sec. 3, we now provide more details about the Spatial-Temporal Sliding Window Attention (ST-SWA).
Previous studies [xi2025sparse] have shown that attention heads in Video Diffusion Transformers exhibit two characteristic sparse patterns: spatial local attention, where a token primarily attends to tokens within the same frame, and temporal attention, where a token focuses on tokens at the same spatial location across neighboring frames. In these models, 3D video latents of size $H_{h},W_{h},F_{h}]$ are flattened into a 1D token sequence.
@@ 0.A.1.1 Spatial-Temporal Sliding Window Attention
In our implementation of spatial-temporal attention, we adopt the headwise permutation kernel from SVG [xi2025sparse] and the headwise sliding-window attention kernel from DiTFastAttnV2 [zhang2025ditfastattnv2]. For configuring the window size of each attention head, we follow the procedure of DiTFastAttnV2 and use eight VBench video prompts as the profiling dataset. We then greedily select the smallest window size whose induced error does not exceed a pre-defined threshold. The error threshold $\delta$ (maximum RSE after applying the sparse attention) is set to $2.0$.
0.A.1.2 Top-K Block Sparse Attention.
@@
Figure 8: Impact of Multi-level Static-Dynamic Scaling on linear attention Regulation and Rank Restoring. Both metrics are evaluated on the attention outputs of a Wan2.1-1.3B model at sampling timestep $t=20$. (Top) The Frobenius norms of attention outputs across different Transformer blocks. The results demonstrate that our multi-level scaling strategy effectively down-weights the linear attention output magnitude to a much smaller range relative to the sparse branch. (Bottom) The effective rank ($r_{\mathrm{eff}}$) across blocks. While naive summation leads to a rank erosion, introducing our scaling strategy progressively restores the joint rank. Finally, the green curve (sparse + linear attention w/ static-dynamic scaling) matches the effective rank of the blue curve (pure sparse attention).
0.A.2 Detailed Analysis of Multi-level Static-Dynamic Scaling
In Sec. 4.1, we identified that there existing rank imbalance and rank degradation (see the Fig. 6) in naive sparse-linear attention. To address these, we introduced a multi-level static-dynamic scaling strategy in Sec. 4.2. In this section, we provide a structured analysis to elucidate how this strategy enables effective sparse and linear attention branches fusion.
Note that, in our analysis, we measure the representation rank using the effective rank metric, which is defined based on the entropy of normalized singular values. Given the singular values $\{\sigma_{i}\}$ of a matrix, we compute
$p_{i}=\frac{\sigma_{i}}{\sum_{j}\sigma_{j}},$ (6)
and define the effective rank as
$r_{\mathrm{eff}}=\exp\Bigg(-\sum_{i}p_{i}\log p_{i}\Bigg).$ (7)
This metric captures the intrinsic dimensionality of the representation and provides a stable estimate of rank under noisy or approximately low-rank conditions.
0.A.2.1 Linear Attention Regulation and Rank Restoration
To regulate the contribution of the linear attention branch, our scaling strategy explicitly controls its numerical magnitude before aggregation. To visualize this effect, we measure the Frobenius norm of the outputs of the linear attention branch, defined as $\|\mathbf{X}\|_{F}=\sqrt{\sum_{i,j}X_{ij}^{2}}$ for a matrix $\mathbf{X}$.
As illustrated in Fig. 8 (Top), the naive linear attention output often exhibits a Frobenius norm close to that of the sparse branch, indicating that the linear branch contributes a comparable magnitude to the final representation. However, as shown in Fig. 8 (Bottom), the two branches exhibit a severe rank imbalance, where the linear attention output is significantly lower-rank than the sparse branch. Directly summing these two branches therefore leads to rank erosion, resulting in a combined representation whose effective rank is even lower than that of the sparse attention alone.
With the introduction of our scaling strategy, the magnitude of the linear attention branch is progressively suppressed. As shown in Fig. 8 (Top), applying the static projection already reduces the energy of the linear branch, while further incorporating the input-timestep-aware dynamic scaling suppresses it even more. Correspondingly, the effective rank of the fused representation gradually recovers, as illustrated in Fig. 8 (Bottom).
Notably, although the layer-wise static projection does not explicitly impose a down-weighting constraint during training, it nevertheless learns to attenuate the contribution of the linear attention branch, implicitly indicating the necessity of regulating linear attention. Furthermore, the dynamic scaling mechanism further reduces the magnitude of the linear branch outputs, validating its complementary role in controlling their influence.
Overall, as the contribution of the linear branch is progressively regulated, the rank of the combined representation increases accordingly, demonstrating that the proposed scaling strategy effectively mitigates the adverse impact of the low-rank linear branch on the sparse attention outputs.
Figure 9: PCA subspace overlap. We collect the output of sparse and linear attention at diffusion timestep 20 across different blocks in WAN1.3B and compute the overlap score in PCA subspace between them.
Figure 10: Relative Singular Value Enhancement.
0.A.2.2 Subspace Decoupling and Spectral Re-attribution
Beyond balancing the contribution of sparse and linear attention branches, we further investigate scaling strategy can help the two branch learn different information.
Subspace Overlap Analysis. To quantify the relationship between the sparse and linear branches, we measure the overlap between their principal subspaces. Specifically, given the output matrices produced by the sparse attention branch $\mathbf{A}$ and the linear attention branch $\mathbf{B}$, we compute their top-$k$ principal components via PCA. Let $\mathbf{V}_{A}$ and $\mathbf{V}_{B}$ denote the corresponding orthonormal bases. We then compute the overlap score
$\text{Overlap}(\mathbf{A},\mathbf{B})=\frac{1}{k}\|\mathbf{V}_{A}^{\top}\mathbf{V}_{B}\|_{F}^{2},$
which measures the average squared cosine similarity between the two $k$-dimensional subspaces.
As shown in Fig. 9, introducing the proposed scaling strategy reduces the overlap score of the two branches. This indicates that the linear branch becomes less aligned with the dominant subspace of the sparse branch. In other words, the scaling mechanism encourages the two branches to encode complementary information rather than redundantly modeling the same dominant directions.
Spectral Tail Revitalization. To further analyze how the linear branch affects the representation spectrum, we perform a singular value decomposition (SVD) analysis on the attention outputs. Let $\sigma_{i}(\cdot)$ denote the $i$-th singular value. We measure the relative singular value change after introducing the linear branch:
$\text{RelSV}_{i}=\frac{\sigma_{i}(\mathbf{A}+\lambda\mathbf{B})-\sigma_{i}(\mathbf{A})}{\sigma_{i}(\mathbf{A})}.$
Interestingly, although the overall effective rank of the representation remains comparable to that of the sparse attention baseline, the singular value distribution changes noticeably after introducing the scaled linear branch. As illustrated in Fig. 10 (Right), the leading singular values and the tail-end singular values both exhibit a increase compared with the sparse attention output. This suggests that the linear branch can enriches the weaker spectral components, encouraging the sparse attention backbone to model additional information that was previously underrepresented.
In contrast, when the two branches are combined via naive summation (without scaling), Fig. 10 (Left) reveals that the dominant singular value of the sparse attention increases dramatically, far more than the rest of the spectrum. This indicates that the linear branch primarily reinforces directions already captured by sparse attention, leading to redundant representations rather than complementary information. Therefore, the proposed static-dynamic scaling strategy plays a crucial role in preventing spectral dominance and enabling the linear branch to act as a complementary refinement to the sparse backbone.
0.A.2.3 Theoretical Interpretation via Weyl’s Inequality
The effectiveness of our scaling strategy in mitigating rank erosion can be qualitatively understood through Weyl’s inequality. For a matrix $\mathbf{A}$ perturbed by another matrix $\mathbf{B}$, the change of each singular value is bounded by the spectral norm (largest singular value) of the perturbation:
$|\sigma_{i}(\mathbf{A}+\mathbf{B})-\sigma_{i}(\mathbf{A})|\leq\|\mathbf{B}\|_{2}.$ (8)
This inequality implies that the magnitude of the perturbation $\mathbf{B}$ directly controls how strongly the singular value spectrum of $\mathbf{A}$ can be altered. In the naive summation setting, the linear branch may have a relatively large norm, allowing it to significantly distort the spectrum of the sparse attention output. In particular, unfavorable interactions between the two branches can suppress weaker singular directions, leading to the observed rank erosion.
Our scaling strategy effectively reduces the overall magnitude of the linear branch before aggregation, thereby tightening the Weyl bound and limiting the possible perturbation to the singular values of $\mathbf{A}$. As a result, the dominant spectral structure of the sparse attention branch is preserved while allowing the linear branch to inject complementary information in less dominant directions.
Our hierarchical scaling transforms the integration into a constrained perturbation: $|\sigma_{i}(\mathbf{A}+\lambda\mathbf{B})-\sigma_{i}(\mathbf{A})|\leq\lambda\|\mathbf{B}\|_{2}$. By enforcing $\lambda\ll 1$, we ensure the perturbation remains small relative to the significant singular values of $\mathbf{A}$. This prevents spectral collapse while allowing the linear branch to provide complementary information.
0.A.3 Details of Post-training Branch Dropping
In Sec. 4.3, we mentioned that some linear attention branches can be dropped to gain more efficiency while maintaining performance. We now provide more details.
We first analyze the distribution of the input-timestep-aware dynamic scaling values across denoising timesteps. The scaling value reflects the relative contribution of the linear attention branch under a specific input prompt and timestep. Since one scaling factor is assigned to each block, it also implicitly captures the block-wise demand for the linear attention branch. Importantly, these scaling values can be computed with negligible overhead during inference, making them a natural indicator for deciding whether a linear attention branch should be executed or skipped. Therefore, we introduce a threshold-based rule: if the scaling value of a branch falls below a threshold, the corresponding linear attention branch is dropped to reduce computation.
As illustrated in Fig. 11, the overall magnitude of scaling values varies across different diffusion timesteps. In particular, the 20th, 40th, 60th, and 80th percentiles consistently shift toward smaller values at later timesteps. In Fig. 12, we further observe that for different input prompts, the quantile statistics at the same timestep are highly consistent, indicating that the distribution of scaling values is largely timestep-dependent rather than prompt-dependent.
Motivated by this observation, we adopt a shared threshold for all blocks at the same timestep. Specifically, for a model with $N$ blocks, dropping the linear attention branches of $M$ blocks corresponds to a drop ratio of $M/N$. By controlling this ratio through a timestep-wise threshold $\tau_{t}$, we can regulate the fraction of linear branches that are skipped at each timestep. Since the scaling statistics are stable across prompts, the thresholds $\tau_{t}$ can be determined using a small calibration set.
We explore two strategies for determining $\tau_{t}$. First, in a coarse-grained strategy, we determine a threshold $\tau_{0}$ at timestep $t=0$ (the input latent is pure noise) such that a desired fraction of linear branches is dropped. We then apply the same threshold to all timesteps: $\tau_{t}=\tau_{0}$ for $t=0,\dots,T-1$, where T is the number of diffusion sampling timesteps.
Because the scaling values naturally vary across timesteps, this constant threshold automatically results in different dropping ratios at different stages of the denoising process, thereby adapting to the varying demand for linear attention.
Second, in a fine-grained strategy, we manually specify the desired dropping ratio for several timestep intervals and calibrate the corresponding thresholds $\tau_{t}$ for each interval using the same calibration set. This provides finer control over the dropping behavior while still requiring only lightweight calibration.
Figure 11: Quantile trends of dynamic scaling values across diffusion timesteps. We plot several quantiles of the input-timestep-aware scaling values over the denoising process. These quantiles reflect the overall magnitude of scaling factors. The consistent downward shift at later timesteps indicates that the contribution of the linear attention branch generally diminishes as the denoising approaches the final image..
Figure 12: Comparison of dynamic scaling values quantiles across timesteps for two different prompts. The quantile curves exhibit highly similar trends between prompts, indicating that the distribution of dynamic scaling values generalizes well across prompts.
We conduct an ablation study on the two dropping schemes using a subset of VBench, as shown in Tab. 6. We find that even the simple coarse-grained strategy achieves performance comparable to SALAD without any dropping, while providing a slight efficiency gain.
In Tab. 2, we also adopt the coarse variant as the drop strategy for SALAD w/ Post-tuning-Drop, achieving an average branch dropping ratio of about $50\%$ across different timesteps.
For comparison, we also evaluate a random dropping strategy that removes the same proportion of linear attention branches. The performance of this baseline is noticeably worse, further demonstrating that the input-timestep-aware dynamic scaling values provide a reliable indicator for determining which linear branches can be safely pruned.
Table 6: Qualitative results with different drop strategies
Drop Strategy SC (↑) BC (↑) IQ (↑) TC (↑) Overall Latency of inference(s)
w.o. Drop 97.21 96.83 69.41 25.56 98.9
fine-grained 97.32 95.84 69.27 25.08 96.6
coarse-grained 97.44 95.78 69.76 24.66 97.4
Random 96.89 96.27 66.79 24.40 96.8
fine-grained drops $20\%$ of branches during the first $40$ timesteps and $30\%$ during the last $10$ timesteps.
coarse-grained removes branches whose dynamic scaling values fall below the lowest $20\%$ at the first timestep, using this threshold uniformly across all timesteps.
Random drops each branch independently with a probability of $20\%$.
0.A.4 Ablation on the Non-Linear Function in the Input-Timestep-Aware Scaling Module
In this section, we investigate the choice of the non-linear function used in the Dynamic Scaler. As illustrated in the main paper Fig. 3, we introduce a non-linearity to increase the expressive capacity of the Input-Timestep-Aware Scaling Module, and we adopt the sigmoid function since a gating range constrained to $[0,1]$ empirically improves the performance of SALAD. Tab. 7 reports the ablation results across different non-linear functions. Sigmoid yields the best performance, while ReLU and Tanh underperform. This suggests that restricting the dynamic scaling values to $[0,1]$ is sufficient and beneficial, which is also consistent with our visualization results.
Table 7: Comparison of Non-Linear Functions in the Dynamic Scaler
Non-Linear Function SC (↑) BC (↑) IQ (↑) TC (↑)
Tanh 96.89 96.25 66.20 26.97
ReLU 96.77 96.10 67.12 28.26